# revtex4-2Repair the float

### v1

@@ IV.1 Special superoperators: complexities
All the algorithms L/R-PTM, M-PTM, C-PTM, and AC-PTM can be briefly summarized as applying the TPD algorithm to the input(s) alongside summation of all fully expanded terms where the single-qubit tensor factors can be drawn from a static lookup-table. Accordingly, the runtime analysis will be quite similar for all these cases and mainly follows the analysis of the worst-case scaling of the TPD [18, Section 4].
Consider first the L/R-PTM for constructing the PTM of single-sided left/right multiplication with some operator $A\in\operatorname{Mat}(2^{n})$. Halving the matrix dimensions in every step, the recursion depth of the algorithm is precisely $n$. At each recursion level $i$, $4^{i}$ CMWs are calculated which each involve taking the sum/difference of two $2^{n-i}\times 2^{n-i}$ matrices, respectively, and the subsequent element-wise division by two. This accounts for $2\cdot 4^{n})$ operations for the mere calculation of the CMWs per step. Additionally, for each of the $4^{i}$ Pauli strings of length $i$ at recursion level $i$, a Kronecker product of the respective elementary PTM (a $4\times 4$-matrix) with the result of the recursively called L/R-PTM has to be calculated and added to the final result $B$. The matrix returned by L/R-PTM is of dimension $4^{n-i}\times 4^{n-i}$, thus calculating any of the Kronecker products at depth $i$ involves $4\cdot 4\cdot 16^{n-i}\cdot 16^{n-i}=16^{n-i+1}$ many operations. The same number of operations is incurred when adding the resulting Kronecker product to $B$, introducing a constant factor of $2$. In total, we obtain
@@ IV.2 Changing channel representation: complexities
algorithm asymptotic complexity
Kraus-PTM $\mathcal{O}(mn16^{n})$
Can-PTM $\mathcal{O}(n16^{n})$
Choi-PTM $\mathcal{O}(n16^{n})$
Chi-PTM $\mathcal{O}(n16^{n})$
IV.3 Special superoperators: numerical experiments
random diagonal matrix random dense matrix
@@ IV.3 Special superoperators: numerical experiments
Figure 5: Measured execution times of L-PTM, M-PTM, C-PTM, and AC-PTM on random matrices and random diagonal matrices. Each plot depicts the respective algorithm’s execution time on a logarithmic axis. Each algorithm admits an exponential scaling with the number of qubits as derived in the complexity analysis and performs significantly better on random diagonal matrices than on general random matrices. The difference in performance is the largest for the M-PTM.
@@ IV.4 Changing channel representation: numerical experiments
random diagonal matrix random matrix Qiskit TPD Hybrid
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$111122223333444455556666777788889999Refer to caption
Execution Time $s$
Number of Qubits $n$Can-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$111122223333444455556666777788889999Refer to caption
Execution Time $s$
Number of Qubits $n$Kraus-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$111122223333444455556666777788889999Refer to caption
Execution Time $s$
Number of Qubits $n$Choi-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$111122223333444455556666777788889999Refer to caption
Execution Time $s$
Number of Qubits $n$Chi-PTM
Figure 6: Measured execution times of Can-PTM, Choi-PTM, Chi-PTM, the respective conversion algorithms in Qiskit (also for Kraus-to-PTM) as well as hybrid version, utilizing Qiskit to translate to the canonical representation and Can-PTM to translate to the respective PTM. The test instances are given by random matrices and random diagonal matrices for systems with up to seven qubits. Each plot depicts the respective algorithms’ execution time on a logarithmic axis. In each case, Qiskit’s implementation starts out with a favorable runtime. On the diagonal test cases, the Can-PTM is able to undercut the former’s runtime from six qubits on; for general random matrices, Can-PTM requires seven qubits to beat Qiskit’ implementation. This behavior is reflected in the favorable runtimes of the hybrid method over the respective conversion algorithm in Qiskit for the other three changes of representations. For up to five qubits, Choi-PTM and Chi-PTM admit better runtimes than the hybrid versions; for six and seven qubits, there runtimes almost match. Each depicted runtime curve is extrapolated for eight and nine qubits by taking the geometric mean of all empirical slopes as the slope of the extrapolation, suggesting favorable runtimes of Can-PTM, Choi-PTM, Chi-PTM and the hybrid versions over Qiskit’s respective conversion algorithm.
@@
For translating different channel representations into the PTM, we compare our proposed algorithms to implementations found in prominent frameworks such as and Qiskit. The results are depicted in Figure 6. We draw exemplary canonical, Choi and Chi matrices from $4^{n}\times 4^{n}$-dimensional random matrices and random diagonal matrices, and $n$ Kraus operators from $2^{n}\times 2^{n}$-dimensional random matrices and random diagonal matrices. Those do not necessarily represent a quantum channel, since we do not check the conditions of complete positivity or trace preservation, but are sufficient to be used a benchmarking instances. For the conversion from Choi/Chi to PTM we test Choi-PTM/Chi-PTM against Qiskit’s implementation and a hybrid version, where we translate the Choi/Chi matrix first to the canonical representation via the suitable Qiskit function and then use Can-PTM to translate the canonical representation to the PTM. When testing Can-PTM itself, the third option is, of course, not applicable. When generating the data, we observed that the runtime of Kraus-PTM massively exceeded the one of Qiskit’s analouge across all qubit numbers. In order to keep the plot range more compact, we omit the graph of Kraus-PTM’s runtimes and only compare Qiskit’s implementation and the hybrid version. Throughout all test instances we observe that Qiskit’s respective conversion algorithm starts out with favorable runtimes, but admits a enormous scaling in the number of qubits. For random diagonal matrices we find that, for all conversions, either our respective algorithm or the proposed hybrid variant starts to perform better than the pure Qiskit implementation, while admitting a more favorable runtime. For general random matrices, the required number of qubits for a speed-up of either our algorithm or the hybrid version is seven, again with an improved empirical scaling. For qubit numbers of eight and nine we could not provide numerical example due to memory issues. Instead, we extrapolate all curves by taking the geometric mean over all empirical slopes as the slope of the extrapolation.
@@ Acknowledgements.
LB acknowledges financial support by the Quantum Valley Lower Saxony. LH acknowledges financial support by the Alexander von Humboldt foundation. We thank Tobias J. Osborne, Luis Santos, Bence Marton Temesi, and Henrik Wilming for helpful discussions. Data and code availability statement. The depicted data is available upon reasonable request from the corresponding author. The code used can be found at https://github.com/HANTLUK/PTM.

### v2

@@ IV.1 Special superoperators: complexities
All the algorithms L/R-PTM, M-PTM, C-PTM, and AC-PTM can be briefly summarized as applying the TPD algorithm to the input(s) alongside summation of all fully expanded terms where the single-qubit tensor factors can be drawn from a static lookup-table. Accordingly, the runtime analysis will be quite similar for all these cases and mainly follows the analysis of the worst-case scaling of the TPD [13, Section 4].
Consider first the L/R-PTM for constructing the PTM of single-sided left/right multiplication with some operator $A\in\operatorname{Mat}(2^{n})$. Halving the matrix dimensions in every step, the recursion depth of the algorithm is precisely $n$. At each recursion level $i$, $4^{i}$ CMWs are calculated which each involve taking the sum/difference of two $2^{n-i}\times 2^{n-i}$ matrices, respectively, and the subsequent element-wise division by two. This accounts for $2\cdot 4^{i}\cdot 4^{n-i}=2\cdot 4^{n}$ operations for the mere calculation of the CMWs per step. Additionally, for each of the $4^{i}$ Pauli strings of length $i$ at recursion level $i$, a Kronecker product of the respective elementary PTM (a $4\times 4$-matrix) with the result of the recursively called L/R-PTM has to be calculated and added to the final result $B$. The matrix returned by L/R-PTM is of dimension $4^{n-i}\times 4^{n-i}$, thus calculating any of the Kronecker products at depth $i$ involves $16\cdot 16^{n-i}=16^{n-i+1}$ many operations. The same number of operations is incurred when adding the resulting Kronecker product to $B$, introducing a constant factor of $2$. In total, we obtain
@@ IV.2 Changing channel representation: complexities
random diagonal matrix random dense matrix Slope
@@
Figure 5: Measured execution times of L-PTM, M-PTM, C-PTM, and AC-PTM on random matrices and random diagonal matrices. Each plot depicts the respective algorithm’s execution time on a logarithmic axis as well as the slope of it’s asymptotic runtime (compare Table 3). Each algorithm admits an exponential scaling with the number of qubits as derived in the complexity analysis and performs significantly better on random diagonal matrices than on general random matrices. The difference in performance is the largest for the M-PTM.
random diagonal matrix random matrix Qiskit TPD Hybrid Slope
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$1111222233334444555566667777Refer to caption
Execution Time $s$
Number of Qubits $n$Can-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$1111222233334444555566667777Refer to caption
Execution Time $s$
Number of Qubits $n$Kraus-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$1111222233334444555566667777Refer to caption
Execution Time $s$
Number of Qubits $n$Choi-PTM
$10^{-4}$$10^{-2}$$10^{0}$$10^{2}$$10^{4}$$10^{6}$1111222233334444555566667777Refer to caption
Execution Time $s$
Number of Qubits $n$Chi-PTM
Figure 6: Measured execution times of Can-PTM, Choi-PTM, Chi-PTM, the respective conversion algorithms in Qiskit (also for Kraus-to-PTM) as well as hybrid version, utilizing Qiskit to translate to the canonical representation and Can-PTM to translate to the respective PTM. The test instances are given by random matrices and random diagonal matrices for systems with up to seven qubits. Each plot depicts the respective algorithms’ execution time as well as the slope of its asymptotic runtime (compare Table 4) on a logarithmic axis. In each case, Qiskit’s implementation starts out with a favorable runtime. On the diagonal test cases, the Can-PTM is able to undercut the former’s runtime from six qubits on; for general random matrices, Can-PTM requires seven qubits to beat Qiskit’ implementation. This behavior is reflected in the favorable runtimes of the hybrid method over the respective conversion algorithm in Qiskit for the other three changes of representations. For up to five qubits, Choi-PTM and Chi-PTM admit better runtimes than the hybrid versions; for six and seven qubits, their runtimes almost match.
algorithm asymptotic complexity
Kraus-PTM $\mathcal{O}(mn16^{n})$
Can-PTM $\mathcal{O}(n16^{n})$
Choi-PTM $\mathcal{O}(n16^{n})$
Chi-PTM $\mathcal{O}(n16^{n})$
IV.3 Special superoperators: numerical experiments
@@ IV.4 Changing channel representation: numerical experiments
For translating different channel representations into the PTM, we compare our proposed algorithms to implementations found in prominent frameworks such as and Qiskit. The results are depicted in Figure 6. We draw exemplary canonical, Choi and Chi matrices from $4^{n}\times 4^{n}$-dimensional random matrices and random diagonal matrices, and $n$ Kraus operators from $2^{n}\times 2^{n}$-dimensional random matrices and random diagonal matrices. Those do not necessarily represent a quantum channel, since we do not check the conditions of complete positivity or trace preservation, but are sufficient to be used a benchmarking instances. For the conversion from Choi/Chi to PTM we test Choi-PTM/Chi-PTM against Qiskit’s implementation and a hybrid version, where we translate the Choi/Chi matrix first to the canonical representation via the suitable Qiskit function and then use Can-PTM to translate the canonical representation to the PTM. When testing Can-PTM itself, the third option is, of course, not applicable. When generating the data, we observed that the runtime of Kraus-PTM massively exceeded the one of Qiskit’s analouge across all qubit numbers. In order to keep the plot range more compact, we omit the graph of Kraus-PTM’s runtimes and only compare Qiskit’s implementation and the hybrid version. Throughout all test instances we observe that Qiskit’s respective conversion algorithm starts out with favorable runtimes, but admits a enormous scaling in the number of qubits. For random diagonal matrices we find that, for all conversions, either our respective algorithm or the proposed hybrid variant starts to perform better than the pure Qiskit implementation, while admitting a more favorable scaling. For general random matrices, the required number of qubits for a speed-up of either our algorithm or the hybrid version is seven, again with an improved empirical scaling.
@@ Acknowledgements.
LB acknowledges financial support by the Quantum Valley Lower Saxony. LH acknowledges financial support by the Alexander von Humboldt foundation. We thank Onur Danaci, Tobias J. Osborne, Luis Santos, Bence Marton Temesi, and Henrik Wilming for helpful discussions. Data and code availability statement. The depicted data as well as the source code are publicly available under [17].